OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned?

OpenStudy (anonymous):

m+w=120 2m-3w=20 m=120-w 2(120-w)-3w=20 240-2w-3w=20 240-5w=20 220=5w w=44 m=76